# Improved Regret of Linear Ensemble Sampling

### v1

@@ Appendix D Proof of Lemma 3
We bound the right hand side using the definition of $\mathcal{E}_{2,t}$.
@@ Appendix E Proof of Lemma 4
holds with probability at least $1-\delta/T$. Note that the right hand side is equal to the definition of $\widetilde{\gamma}_{T}$, defined in Eq. (7). ∎
Appendix F Rigorous Justification of Claim 1
In this section, we rigorously justify Claim 1 that is stated in the proof of Theorem 1. To do so, we present a different viewpoint on the perturbation sequences.
@@ Appendix F Rigorous Justification of Claim 1
$\text{rep}(v,a,n):=\begin{pmatrix}v&\ldots&v&\boldsymbol{0}_{d}&\ldots&\boldsymbol{0}_{d}\end{pmatrix}\in\mathbb{R}^{d\times n}\,.$
@@
$\mathbf{\Phi}_{t}=\begin{pmatrix}\sqrt{\lambda}I_{d}&\text{rep}(x^{1},N_{1,t},T)&\ldots&\text{rep}(x^{K},N_{K,t},T)\end{pmatrix}\in\mathbb{R}^{d\times(d+KT)}\,.$
@@
We apply Azuma-Hoeffding inequality to show that $\frac{1}{m}\sum_{j=1}^{m}I^{j}\left(\mathbf{\Phi}_{t-1}\right)$ is bounded below with high probability. Since $\{I^{j}(\mathbf{\Phi}_{t-1})\}_{j=1}^{m}$ are i.i.d. Bernoulli random variables with the associated probability greater than $p_{N}/2$, it holds that
@@ Appendix H Generalizability of Perturbation Distributions
holds with probability at least $1-2\delta$, which proves the the lemma. ∎
We also provide that the $\ell_{2}$-norm of the vector $W$ whose components are i.i.d. samples of a subGaussian distribution is upper-bounded with high-probability. This lemma, combined with Lemma 9, justifies $\mathcal{P}_{I}$ to be a distribution over $\mathbb{R}^{d}$ such that each component is an i.i.d. sample of a subGaussian distribution.

### v3

@@ Appendix D Proof of Lemma 3
We bound the right-hand side using the definition of $\mathcal{E}_{2,t}$.
@@ Appendix E Proof of Lemma 4
holds with probability at least $1-\delta/T$. Note that the right-hand side is equal to the definition of $\widetilde{\gamma}_{T}$, defined in Eq. (7). ∎
Appendix F Rigorous Justification of Proposition 1
In this section, we rigorously justify Proposition 1 that is stated in the proof of Theorem 1. To do so, we present a different viewpoint on the perturbation sequences.
@@ Appendix F Rigorous Justification of Proposition 1
$\mathrm{rep}(v,a,n):=\begin{pmatrix}v&\ldots&v&\boldsymbol{0}_{d}&\ldots&\boldsymbol{0}_{d}\end{pmatrix}\in\mathbb{R}^{d\times n}\,.$
@@
$\mathbf{\Phi}_{t}=\begin{pmatrix}\sqrt{\lambda}I_{d}&\mathrm{rep}(x^{1},N_{1,t},T)&\ldots&\mathrm{rep}(x^{K},N_{K,t},T)\end{pmatrix}\in\mathbb{R}^{d\times(d+KT)}\,.$
@@
We apply the Azuma-Hoeffding inequality to show that $\frac{1}{m}\sum_{j=1}^{m}I^{j}\left(\mathbf{\Phi}_{t-1}\right)$ is bounded below with high probability. Since $\{I^{j}(\mathbf{\Phi}_{t-1})\}_{j=1}^{m}$ are i.i.d. Bernoulli random variables with the associated probability greater than $p_{N}/2$, it holds that
@@ Appendix H Generalizability of Perturbation Distributions
holds with probability at least $1-2\delta$, which proves the lemma. ∎
We also provide that the $\ell_{2}$-norm of the vector $W$ whose components are i.i.d. samples of a subGaussian distribution is upper-bounded with high probability. This lemma, combined with Lemma 9, justifies $\mathcal{P}_{I}$ to be a distribution over $\mathbb{R}^{d}$ such that each component is an i.i.d. sample of a subGaussian distribution.